An elf steals a sleeping dragon's gold treasure and runs off at 16 mph. 3 hours after the theft, the dragon wakes up and sets off after the elf at 20 mph. How long will it take the dragon to capture his lunch?
Dragon dines on firey roasted elf in ________ hours.

Answer by stanbon(75887)

An elf steals a sleeping dragon's gold treasure and runs off at 16 mph.
---------------------------------
3 hours after the theft, the dragon wakes up and sets off after the elf at 20 mph.
-------------------------------------------
How long will it take the dragon to capture his lunch?
------------------------------------
Elf DATA:
rate = 16 mph ; time = x hrs ; distance = 16x miles
------
Dragon DATA:
rate = 20 mph ; time = x-3 hrs ; distance = 20(x-3) miles
----------------------
Equation:
elf distance = dragon distance
16x = 20(x-3)
16x = 20x - 60
4x = 60
x = 15 hrs (elf time)
x-3 = 12 hrs (dragon time)
==============================
Cheers,
Stan H.
================




Dragon dines on firey roasted elf in ________ hours.